xxsaucegodxx:

A ride at an amusement park moves the riders in a circle at a rate of 6.0 m/s. If the radius of the ride is 9.0 meters, what is the acceleration of the ride? 4.0 m/s2 0.67 m/s2 0.075 m/s2 54 m/s2

Mercury:

centripetal acceleration: a = v^2 / r plug in your velocity (v = 6.0 m/s) and your radius (r = 9.0m) and evaluate a.